# Machine Learning for Synthetic Data Generation: A Review

### v8

@@ Abstract
data synthesis, machine learning, data trustworthiness
@@ I Introduction
  • We outline several general evaluation strategies to assess the quality of synthetic data (Section VIII).
  • We identify challenges faced in generating synthetic data and during the deployment process, highlighting potential future work that could further enhance functionality (Section IX).
@@ II-E Business
Efforts have been made to secure original data using technologies like encryption, anonymization, and cutting-edge privacy preservation [101]. However, information gleaned from the data may still be employed to trace individuals, thereby posing the risk [102]. A notable advantage of synthetic data lies in its ability to eliminate the exposure of critical data, thus ensuring privacy and security for both companies and their customers [103].
Moreover, synthetic data enables organizations to access data more rapidly, as it bypasses privacy and security protocols [104]. In the past, institutions possessing extensive data repositories could potentially assist decision-makers in resolving a broad spectrum of issues. However, accessing such data, even for internal purposes, was hindered by confidentiality concerns. Presently, companies are harnessing synthetic data to refresh and model original data, generating continuous insights that contribute to enhancing the organization’s performance [4].
@@ II-F Education
Furthermore, synthetic data can be a powerful tool in teacher training programs. For example, teacher candidates can use synthetic student data to practice data-driven instructional strategies, including differentiated instruction and personalized learning plans. They can analyze this synthetic data, identify patterns, determining student needs, and adjusting their instructional plans accordingly. By using synthetic data, teacher candidates gain practical experience in analyzing student data and adapting their teaching without infringing on the privacy of actual students [107]. Thus, synthetic data serves as a valuable bridge between theory and practice in education, driving innovation while safeguarding privacy.
@@ II-G Location and Trajectory Generation
Location and trajectory are a particular form of data that could highly reflect users’ daily lives, habits, home addresses, workplaces, etc. To protect location privacy, synthetic location generation is introduced as opposed to location perturbation [108]. The main challenge of generating synthetic location and trajectory data is to resemble genuine user-produced data while offering practical privacy protection simultaneously. One approach to generating the location and trajectory data is to inject a synthetic point-based site within a user’s trajectory [109, 110, 111]. Synthetic trajectory generation is frequently combined with privacy-enhancing techniques to further prevent sensitive inference from the synthesized data. For example, Chen et al. [112] introduces an N-gram-based method to predict the following position based on previous positions for publishing trajectory. They exploit the prefix tree to describe the n-gram model while combining it with differential privacy [113]. [114] proposes a synthetic trajectory strategy based on the discretization of raw trajectories using hierarchical reference systems to capture individual movements at differing speeds. Their method adaptively selects a small set of reference systems and constructs prefix tree counts with differential privacy. Applying direction-weighted sampling, the decrease in tree nodes reduces the amount of added noise and improves the utility of the synthetic data. By extracting multiple differential private distributions with redundant information, [115] the author generated a new trajectory with samples from these distributions. In addition to differential privacy, Bindschaedler and Shokri [116] enforce plausible deniability to generate privacy-preserving synthetic traces. It first introduces trace similarity and intersection functions that map a fake trace to a real hint under similarity and intersection constraints. Then, it generates one fake trace by clustering the locations and replacing the trajectory locations with those from the same group. If the fake trace satisfies plausible deniability, i.e., there exist k other real traces that can map to the fake trace, then it preserves the privacy of the seed trace. While existing studies mainly use the Markov chain model, [117] proposes to choose between the first-order and second-order Markov model adaptively. The proposed PrivTrace controls the space and time overhead by the first-order Markov chain model and achieves good accuracy for next-step prediction by the second-order Markov chain model.
@@ II-H AI-Generated Content (AIGC)
AI-Generated Content (AIGC) stands at the forefront of the technology and content creation industry, changing the dynamics of content production. A typical example of AIGC is OpenAI’s ChatGPT, an AI-driven platform generating human-like text in response to prompts or questions. It leverages a vast corpus of internet text to generate detailed responses, often indistinguishable from those a human writer would produce. This capacity extends beyond simple question-answer pairs to crafting whole articles, stories, or technical explanations on a wide range of topics, thus creating a novel way of producing blog posts, articles, social media content, and more [118, 119].
@@
An innovative application of AIGC is in the field of news reporting. News agencies are increasingly using AI systems, such as the GPT series, to generate news content. For instance, the Associated Press uses AI to generate news articles about corporate earnings automatically. The AI takes structured data about company earnings and transforms it into a brief, coherent, and accurate news report. This automation allows the agency to cover a much larger number of companies than would otherwise be possible with human journalists alone [121].
@@
Additionally, AIGC has found its place in the creative domain, with AI systems being used to generate book descriptions, plot outlines, and even full chapters of novels. For instance, a novelist could use ChatGPT to generate a synopsis for their upcoming book based on a few keywords or prompts related to the story. Similarly, marketing teams utilize AI to create compelling product descriptions for online marketplaces [122]. This not only increases efficiency but also provides a level of uniformity and scalability that would be challenging to achieve with human writers alone. Through these examples, it is clear that AIGC is profoundly impacting the landscape of content creation and will continue to shape it in the future [120].
@@ III-A Multiple Layer Perception (MLP)
Multiple layer perception (MLP) is a classical (or vanilla) feedforward artificial neural network that uses a fully-connected connection in a single layer. It models all the possible interactions between the features. It is also the most used neural network. In the $i$-th layer, the propagation can be written as
@@ III-C Recurrent Neural Network (RNN)
The sequence is one of the most popular data formats in real-world applications, e.g., natural language, speech signal, etc. Recurrent neural network (RNN) was designed to represent and generate sequence data [125]. Suppose we have a sequence of length $T$, $X=[x_{1},x_{2},...,x_{T}]$, $x_{t}$ denotes the $t$-th token, recurrent neural network designs a recursive (or recurrent) mechanism to reuse the neural network module. Specifically, at the $t$-th step, we have
@@ III-D Graph Neural Network (GNN)
There are many graph-structured data in downstream applications, such as social networks, brain networks, chemical molecules, knowledge graphs, etc. Graph neural network was proposed to model graph-structured data and learn the topological structure from graph [128, 129, 130]. Specifically, graph neural networks build the interaction between the connected nodes and edges to model the topological structure of the graph [131, 132, 133, 134]. The feedforward rule of graph neural network can be formulated as
@@
where $h_{i}^{(l)}$ denotes the embedding of the $i$-th node at the $l$-th layer, $m_{j}$ denotes the node feature of the node $j$, $m_{ji}$ denotes the edge feature of the edge that connects nodes j and i, $\mathcal{N}(i)$ is the set of nodes that connect with $i$. Within each layer, we update the current representation via aggregating the information from (1) representation from the previous layer, (2) node feature (3) edge feature. For example, for chemical compounds, each node corresponds to an atom, the node feature is the category of the atom, e.g., Carbon atom, Nitrogen atom, Oxygen atom; each edge is a chemical bond, the edge feature is the type of the bond, including single bond, double bond, triple bond and aromatic bond.
@@ IV-B Self-Supervised Learning (SSL)
Labeled data are expensive to acquire so the number of available labeled data is usually limited. To address this issue, self-supervised learning (SSL) was proposed, which is a learning paradigm that curates the supervision signal from the data itself. It is parallel to supervised learning and unsupervised learning. Different from supervised learning, self-supervised learning can learn from massive unlabeled data. Self-supervised learning is usually used as a pretraining strategy to learn the representation from massive unlabelled data [143]. The core idea of self-supervised learning is to mask a subset of the raw data feature and build a machine learning model to predict the masked data. then the pre-trained machine learning model (usually a neural network) is used as a “warm start”, and is furtherly finetuned for the downstream applications.
@@ IV-D Generative Adversarial Network (GAN)
Generative adversarial network (GAN) [138, 146, 147] formulates the generation problem into a supervised learning task. Specifically, it comprises two neural network modules: discriminator and generator. The objective of the generator is to generate data that are close to the real data, On the other hand, the objective of the discriminator is to discriminate the fake data (generated by the generator) from the real ones. It performs a binary classification task, where the real data from the training set are regarded as the positive samples; the generated data (by generator) are regarded as the negative samples. generator and discriminator are trained in a mini-max manner.
@@
  • optimize discriminator and fix generator: the objective function reduces to a binary classification problem, that is,
@@ IV-E Reinforcement Learning (RL)
The objective function is not differentiable with parameter $\theta$. We use policy gradient to obtain an unbiased estimator of the objective gradient $\nabla_{\theta}\mathcal{L}(\theta)$ [141] and then use stochastic optimization methods to maximize the expected reward. Generating synthesis data can be viewed as sequential decision-making by sequentially generating one basic component (basic structure).
@@ V Privacy Risks and Prevention
Open release and free data exchange would benefit research and industry development. However, there are cases where datasets exist but cannot be publicly disclosed due to privacy concerns. In the meantime, the emergence and the enforcement of privacy regulations such as HIPAA [^1] https://www.hhs.gov/hipaa/index.html, EU General Data Protection Regulation (GDPR [^2] https://gdpr-info.eu) and the California Consumer Privacy Act (CCPA [^3] https://oag.ca.gov/privacy/ccpa) make it infeasible for sensitive data with privacy implications. Regulated data such as clinical and genomics data in raw may not be shared and one solution is to share synthesized data instead.
@@ V-A Privacy Risks in Data Synthesis
Due to the utility goal of data synthesis, the synthesized data tend to preserve the distribution of the original data. Therefore, the deployment upon these models could be subject to privacy leakage. For deep neural network based approaches, membership inference attack [155, 156] would identify if an input is in the training data or not and thus can be used to determine how close the synthesized data is to the original data. At feature level, sensitive attributes such as skin color can be inferred from the behavior of the deep learning model [157] and even the single training instance can be reconstructed [158, 159, 160]. For generative AI models, the generative learning process and the high complexity of the model jointly encourage a distribution that is concentrated around training samples. By repeatedly sampling from the distribution, there is a considerable chance of recovering the training samples [161, 162, 163] or the membership of the training data [164].
@@
[167] differential privacy GAN CNN image
[49] differential privacy GAN Private Aggregation of Teacher Ensembles attribute
[168] differential privacy GAN CNN EHR
[169] differential privacy GAN CNN image and EHR
[14] differential privacy autoencoder autoencoder attribute
[170] differential privacy MRF NA (w.o. DNN) attribute
[171] compressive privacy GAN CNN image
[172] differential privacy GAN CNN image
@@
[175] differential privacy GAN CNN image
[176] differential privacy Bayesian network NA (w.o. DNN) attribute

### v9

@@ Abstract
data synthesis, machine learning, generative modeling
@@ I Introduction
  • We outline several general evaluation strategies to assess the quality of synthetic data (Section VII).
  • We identify challenges faced in generating synthetic data and during the deployment process, highlighting potential future work that could further enhance functionality (Section VIII).
@@ II-E Business
Efforts have been made to secure original data using technologies like encryption, anonymization, and cutting-edge privacy preservation [101]. However, information gleaned from the data may still be employed to trace individuals, thereby posing the risk [102]. A notable advantage of synthetic data lies in its ability to eliminate the exposure of critical data, thus ensuring privacy and security for both companies and their customers [103]. Moreover, synthetic data enables organizations to access data more rapidly, as it bypasses privacy and security protocols [104]. In the past, institutions possessing extensive data repositories could potentially assist decision-makers in resolving a broad spectrum of issues. However, accessing such data, even for internal purposes, was hindered by confidentiality concerns. Presently, companies are harnessing synthetic data to refresh and model original data, generating continuous insights that contribute to enhancing the organization’s performance [4].
@@ II-F Education
Furthermore, synthetic data can be a powerful tool in teacher training programs. For example, teacher candidates can use synthetic student data to practice data-driven instructional strategies, including differentiated instruction and personalized learning plans. They can analyze this synthetic data, identify patterns, determine student needs, and adjust their instructional plans accordingly. By using synthetic data, teacher candidates gain practical experience in analyzing student data and adapting their teaching without infringing on the privacy of actual students [107]. Thus, synthetic data serves as a valuable bridge between theory and practice in education, driving innovation while safeguarding privacy.
@@ II-G Location and Trajectory Generation
Location and trajectory are a particular form of data that could highly reflect users’ daily lives, habits, home addresses, workplaces, etc. To protect location privacy, synthetic location generation is introduced as opposed to location perturbation [108]. The main challenge of generating synthetic location and trajectory data is to resemble genuine user-produced data while offering practical privacy protection simultaneously. One approach to generating the location and trajectory data is to inject a synthetic point-based site within a user’s trajectory [109, 110].
Synthetic trajectory generation is frequently combined with privacy-enhancing techniques to further prevent sensitive inference from the synthesized data. For example, Chen et al. [111] introduces an N-gram-based method to predict the following position based on previous positions for publishing trajectory. They exploit the prefix tree to describe the n-gram model while combining it with differential privacy [112]. [113] extends the n-gram model with local differential privacy and [114] further replaces the n-gram model with key movement mobility for differentially private trajectory generation. By comparison, [115] proposes a synthetic trajectory strategy based on the discretization of raw trajectories using hierarchical reference systems to capture individual movements at differing speeds. Their method adaptively selects a small set of reference systems and constructs prefix tree counts with differential privacy. Applying direction-weighted sampling, the decrease in tree nodes reduces the amount of added noise and improves the utility of the synthetic data. [116] constructs the differentially private prefix tree and calibrates original trajectories against a selection of anchor points. By extracting multiple differential private distributions with redundant information [117, 118], the authors generate a new trajectory with samples from these distributions. By comparison, [119] estimates various distributions of an attribute set to determine trajectories and [120] consider the interactions between different attributes by grouping strongly correlated attributes into non-disjoint sets and constructing a corresponding distribution for each set.
In addition to differential privacy, Bindschaedler and Shokri [121] enforce plausible deniability to generate privacy-preserving synthetic traces. It first introduces trace similarity and intersection functions that map a fake trace to a real hint under similarity and intersection constraints. Then, it generates one fake trace by clustering the locations and replacing the trajectory locations with those from the same group. If the fake trace satisfies plausible deniability, i.e., there exist k other real traces that can map to the fake trace, then it preserves the privacy of the seed trace. While existing studies mainly use the Markov chain model, [122] proposes PrivTrace, which controls the space and time overhead by the first-order Markov chain model and achieves good accuracy for next-step prediction by the second-order Markov chain model. [123] considers the location synthesizer that generates location traces, including co-locations of friends, while offering node-level differential privacy for the friendship and user-level differential privacy for the co-location count matrix.
@@ II-H AI-Generated Content (AIGC)
AI-Generated Content (AIGC) stands at the forefront of the technology and content creation industry, changing the dynamics of content production. A typical example of AIGC is OpenAI’s ChatGPT, an AI-driven platform generating human-like text in response to prompts or questions. It leverages a vast corpus of internet text to generate detailed responses, often indistinguishable from those a human writer would produce. This capacity extends beyond simple question-answer pairs to crafting whole articles, stories, or technical explanations on a wide range of topics, thus creating a novel way of producing blog posts, articles, social media content, etc [124, 125].
@@
An innovative application of AIGC is in the field of news reporting. News agencies increasingly use AI systems, such as the GPT series, to generate news content. For instance, the Associated Press uses AI to generate news articles about corporate earnings automatically. The AI takes structured data about company earnings and transforms it into a brief, coherent, and accurate news report. This automation allows the agency to cover many companies that would be possible with human journalists alone [127].
@@
Additionally, AIGC has found its place in the creative domain, with AI systems being used to generate book descriptions, plot outlines, and even full chapters of novels. For instance, a novelist could use ChatGPT to generate a synopsis for their upcoming book based on a few keywords or prompts related to the story. Similarly, marketing teams utilize AI to create compelling product descriptions for online marketplaces [128]. This increases efficiency and provides a level of uniformity and scalability that would be challenging to achieve with human writers alone. Looking forward, AIGC is profoundly impacting the landscape of content creation and will continue to shape it in the future [126].
II-I Finance
Synthetic data generation offers significant benefits for the finance industry [102], as detailed below. First, financial data is highly sensitive and subject to stringent privacy regulations [129]. Synthetic data mimics real data without exposing actual customer information, enabling institutions to comply with privacy laws while still utilizing detailed datasets for analysis and development. Second, synthetic data can be used to test and validate financial algorithms and models under various conditions. For example, trading algorithms can be tested using synthetic market data to evaluate their performance under different market scenarios, including rare or extreme events that may not be present in historical data [130]. Third, developing and testing financial algorithms requires large volumes of high-quality data. Synthetic data provides an endless supply of training data, enabling thorough backtesting of trading strategies and machine learning models without the risk of overfitting historical data [131].
Synthetic data generation also transforms the financial services industry by enabling more accurate risk assessments and fraud detection[102]. Synthetic data generation can identify anomalies and potential risks by simulating financial transactions and market behaviors, allowing financial institutions to implement more effective fraud prevention measures and develop more resilient financial strategies. Furthermore, synthetic data generation can support compliance with regulatory requirements by providing detailed, real-time reporting and analysis of financial activities [47]. In the context of human resources, synthetic data generations can model workforce dynamics, including employee performance, engagement, and turnover. By analyzing these models, businesses can develop strategies to improve employee satisfaction, enhance productivity, and reduce turnover rates. For example, synthetic data generations can simulate the impact of various HR policies on workforce morale and performance, helping HR departments to implement the most effective practices.
II-J Other Applications
The techniques for synthetic data generation described in this paper have far-reaching implications beyond the specific domains covered. Here are some notable applications:
  • Retail and Marketing: In retail, synthetic data can model customer interactions, purchasing behaviors, and inventory management [132]. This aids in developing personalized marketing strategies, optimizing supply chains, and improving customer service without infringing on individual privacy.
  • Environmental Studies: Synthetic data can simulate environmental conditions, weather patterns, and ecological interactions [133]. This is particularly useful for studying climate change, biodiversity, and conservation efforts, allowing researchers to test hypotheses and model practical scenarios without the constraints of limited real-world data.
  • Urban Planning and Development: In urban planning, synthetic data can be used to simulate population growth, traffic flows, and infrastructure development [134]. This helps city planners and developers make informed decisions about resource allocation, transportation systems, and sustainable development initiatives.
  • Software Development and Testing: In software development, synthetic code generation can simulate various coding scenarios, bug patterns, and software behaviors [135]. This is particularly useful for testing and debugging, as it allows developers to identify and fix potential issues without the constraints of existing codebases. Synthetic code can also aid in developing personalized coding assistants, optimizing software performance, and improving the reliability of code releases [136]. Additionally, by generating diverse and extensive code samples, developers can enhance machine learning models for code completion and error detection, ultimately leading to more efficient and robust software development processes.
@@ III-A Multiple Layer Perception (MLP)
Multiple-layer perception (MLP) is a classic feedforward artificial neural network that uses a fully connected connection in a single layer. It models all the possible interactions between the features. It is also the most used neural network. In the $i$-th layer, the propagation can be written as
@@ III-C Recurrent Neural Network (RNN)
The sequence is one of the most popular data formats in real-world applications, such as natural language and speech signals. Recurrent neural network (RNN) was designed to represent and generate sequence data [139]. Suppose we have a sequence of length $T$, $X=[x_{1},x_{2},...,x_{T}]$, $x_{t}$ denotes the $t$-th token, recurrent neural network designs a recursive (or recurrent) mechanism to reuse the neural network module. Specifically, at the $t$-th step, we have
@@ III-D Graph Neural Network (GNN)
Many graph-structured data exist in downstream applications, such as social networks, brain networks, chemical molecules, and knowledge graphs. Graph neural network was proposed to model graph-structured data and learn the topological structure from graph [142, 143, 144]. Specifically, graph neural networks build the interaction between the connected nodes and edges to model the topological structure of the graph [145, 146]. The feedforward rule of graph neural network can be formulated as
@@
where $h_{i}^{(l)}$ denotes the embedding of the $i$-th node at the $l$-th layer, $m_{j}$ denotes the node feature of the node $j$, $m_{ji}$ denotes the edge feature of the edge that connects nodes $j$ and $i$, $\mathcal{N}(i)$ is the set of nodes that connect with $i$. Within each layer, we update the current representation via aggregating the information from (1) representation from the previous layer, (2) node feature (3) edge feature. For example, in chemical compounds, each node corresponds to an atom, and the node feature is the category of the atom, e.g., Carbon atom, Nitrogen atom, Oxygen atom; each edge is a chemical bond, and the edge feature is the type of the bond, including single bond, double bond, triple bond, and aromatic bond.
@@ IV-B Self-Supervised Learning (SSL)
Labeled data are expensive to acquire so the number of available labeled data is usually limited. To address this issue, self-supervised learning (SSL) was proposed. This learning paradigm curates the supervision signal from the data itself. It is parallel to supervised learning and unsupervised learning. Different from supervised learning, self-supervised learning can learn from massive unlabeled data. Self-supervised learning is usually used as a pretraining strategy to learn the representation from massive unlabelled data [155]. The core idea of self-supervised learning is to mask a subset of the raw data feature and build a machine learning model to predict the masked data. then the pre-trained machine learning model (usually a neural network) is used as a “warm start”, and is furtherly finetuned for the downstream applications.
@@ IV-D Generative Adversarial Network (GAN)
Generative adversarial network (GAN) [150, 158, 159] formulates the generation problem into a supervised learning task. Specifically, it comprises two neural network modules: discriminator and generator. The objective of the generator is to generate data that are close to the real data, By comparison, the objective of the discriminator is to discriminate the fake data (generated by the generator) from the real ones. It performs a binary classification task, where the real data from the training set are regarded as the positive samples; the generated data (by generator) are regarded as the negative samples. generator and discriminator are trained in a mini-max manner.
@@
  • optimize discriminator and fix generator: the objective function reduces to a binary classification problem,
@@ IV-E Reinforcement Learning (RL)
The objective function is not differentiable with parameter $\theta$. We use policy gradient to obtain an unbiased estimator of the objective gradient $\nabla_{\theta}\mathcal{L}(\theta)$ [153] and then use stochastic optimization methods to maximize the expected reward. Generating synthesis data can be viewed as sequential decision-making by sequentially generating one basic structure.
@@ IV-F Diffusion Model
IV-G Multimodal Learning
Multimodal data refers to datasets that integrate multiple types of data, such as text, images, audio, and numerical values. This type of data provides a comprehensive view by combining different sources of information, which is crucial for tasks requiring a holistic understanding of complex scenarios. In fields like healthcare [167], finance, and autonomous systems, multimodal data enables more accurate and robust analysis and decision-making by leveraging the strengths of each data type. For instance, in drug discovery, multimodal data can combine genomic data, chemical structures, and clinical outcomes to enhance the prediction of drug efficacy and safety [93].
Synthetic multimodal data generation involves creating artificial datasets that integrate multiple types of data, such as text, images, audio, and numerical data, to simulate real-world scenarios. This technique is particularly valuable in fields like healthcare [16], finance [102], and education systems [107], where data is often complex and heterogeneous.
Then, we review some cutting-edge techniques for synthetic multimodal data generation. GANs can generate one type of data from another, such as images from textual descriptions or audio from images. This cross-modal generation capability is essential for creating cohesive multimodal datasets [11]. Recently, ChatGPT [127] supports multimodal data generation, including image, text, and numerical features.
@@ V Privacy Risks and Prevention
Open release and free data exchange would benefit research and industry development. However, there are cases where datasets exist but cannot be publicly disclosed due to privacy concerns. Regulated data, such as clinical and genomics data in raw form, may not be shared, and one solution is to share synthesized data instead.
@@ V-A Privacy Risks in Data Synthesis
Due to the utility goal of data synthesis, the synthesized data tends to preserve the distribution of the original data. Therefore, the deployment of these models could be subject to privacy leakage. For deep neural network-based approaches, membership inference attack [168, 169] would identify if an input is in the training data or not and thus can be used to determine how close the synthesized data is to the original data. At the feature level, sensitive attributes such as skin color can be inferred from the behavior of the deep learning model [170], and even the single training instance can be reconstructed [171, 172, 173]. For generative AI models, the generative learning process and the high complexity of the model jointly encourage a distribution that is concentrated around training samples. By repeatedly sampling from the distribution, there is a considerable chance of recovering the training samples or attributes [174, 175, 176, 177, 178, 179, 180], or the membership of the training data [181].
@@
[14] differential privacy autoencoder autoencoder attribute
[27] differential privacy VAE + GAN recurrent autoencoder EHR
[49] differential privacy (PATE) GAN DNN attribute
[111] differential privacy n-gram NA (w.o. DNN) sequential/time series
[113] local differential privacy n-gram NA (w.o. DNN) trajectory
[114] local differential privacy Markov probabilistic model NA (w.o. DNN) trajectory
[115] differential privacy Markov probabilistic model NA (w.o. DNN) trajectory
[116] differential privacy Markov probabilistic model NA (w.o. DNN) social media trajectory
[118] differential privacy Markov probabilistic model NA (w.o. DNN) trajectory
[119] differential privacy distribution estimation NA (w.o. DNN) location
[120] differential privacy distribution estimation NA (w.o. DNN) trajectory
[121] plausible deniability Hidden Markov Models NA (w.o. DNN) trajectory
[122] differential privacy Markov chain model NA (w.o. DNN) trajectory
[123] differential privacy probabilistical transform NA (w.o. DNN) trajectory
@@
[183] compressive privacy GAN DNN image
[184] differential privacy GAN DNN image
[185] differential privacy GAN DNN image and EHR
[186] differential privacy GAN DNN image
[187] differential privacy GAN DNN image
[188] differential privacy GAN DNN attribute (tabular) and graph
[189] differential privacy GAN DNN image and EHR
[190] differential privacy GAN DNN time series
[191] differential privacy GAN DNN attribute
[192] differential privacy GAN DNN attribute (tabular)
[193] differential privacy GAN DNN image
[194] differential privacy GAN DNN image
[195] differential privacy GAN autoencoder attribute
[196] differential privacy GAN DNN trajectory
[197] differential privacy (PATE) GAN DNN image
[198] local differential privacy GAN DNN spatial point
[199] differential privacy Maximum Mean Discrepancy Hermite polynomial features attribute (tabular) and image
[200] differential privacy Maximum Mean Discrepancy Random feature mean embeddings image
[201] differential privacy Markov random field NA (w.o. DNN) attribute
[202] differential privacy Markov random field NA (w.o. DNN) attribute
[203] differential privacy Probabilistic graphical models NA (w.o. DNN) attribute
[204] differential privacy Maximum Cardinality Matching NA (w.o. DNN) attribute (tabular)
[205] differential privacy Bayesian network NA (w.o. DNN) attribute
[206] differential privacy Bayesian network NA (w.o. DNN) attribute
[208] differential privacy statistical queries A (w.o. DNN) attribute
[209] differential privacy statistical queries A (w.o. DNN) attribute
[212] differential privacy autoencoder autoencoder text/image QA and attribute
@@
[214] differential privacy Langevin Markov chain Monte Carlo Energy-based Model image